110. A U.S. corporation has purchased currency put options to hedge a 100,000 Canadian dollar (C$) receivable. The premium is $.02 per unit and the exercise price of the option is $.94. If the spot rate at the time of maturity is $.99, what is the net amount received by the corporation if it acts rationally?

a. $92,000.

b. $97,000.

c. $96,000.

d. $94,000.

111. A U.S. corporation has purchased currency call options to hedge a ₤70,000 payable. The premium is $.02 [.05] and the exercise price of the option is $.50 [$1.50]. If the spot rate at the time of maturity is $.65 [$1.65], what is the total amount paid by the corporation if it acts rationally?

a. $33,600

b. $46,900

c. $44,100

d. $36,400
Answer #1

A.

Receivable amount=   100000   canadian dollar
Excercise price of put option =$   0.94  
Premium paid=   0.02  
Spot rate at maturity =$   0.99  
Value of put option = Excercise price-Spot rate      
Net profit or loss = Value of put option - Premium paid      
If Spot rate is more than Excercise price, put option will not be excercised. Value shall be nil. in this case loss will be equal to premium paid.       
VP =0.94-0.99=   0   Option will not be excercised
Net profit =0-0.02=   -0.02  
      
Total amount received for receivables = receivables amount *(Spot rate +profit-loss)      
100000*(0.99-0.02)      
97000      
      
So a rationale corporation will not excercise option and Net amount receivables is   $97,000.00  
      

B.

Payable amount=    70000   £70,000
Excercise price of option =$   0.5  
Premium paid=    0.02  
Spot rate at maturity =$   0.65  


Value of call option = Spot rate - Excercise price      
Net profit or loss = Value of call option - Premium paid


If Spot rate is less than Excercise price, call option will not be excercised. Value shall be nil. in this case loss will be equal to premium paid.       
VC = 0.65-0.50=   0.15  
Net profit =0.15-0.02=   0.13  
      
Total amount paid for payables = Payable amount *(Spot rate - Profit)      
70000*(0.65-0.13)      
36400      
      
So a rationale corporation will excercise option and Net amount payable is    $36,400.00  
      

Note : For payable amount, profit is deducted and loss is added. As these are cost.

For receivables amount, profit is added and loss is deducted. As these are revenue

Please thumbs up or post. Comment if any query